# SwiftEmbed: Ultra-Fast Text Embeddings via Static Token Lookup for Real-Time Applications

### v1

@@ 1 Introduction
Text embeddings have become fundamental to natural language processing applications, yet transformer-based approaches introduce prohibitive latency for real-time scenarios requiring sub-millisecond response times. While models like BERT [3] achieve state-of-the-art semantic quality, their multi-layer attention mechanisms create computational bottlenecks that limit deployment in latency-sensitive environments.
@@ 2.1 Transformer-Based Embedding Models
Transformer architectures from BERT [3] to Sentence-BERT [9] have established contemporary standards for semantic representations through multi-layer attention mechanisms [13]. While achieving superior quality, their computational complexity renders them unsuitable for ultra-low latency scenarios. Compression techniques like DistilBERT [11] and TinyBERT [5] reduce model sizes but retain full transformer inference, fundamentally limiting latency improvements. Recent advances in efficient transformers [21] and model compression [20] have explored various optimization paths while maintaining transformer architecture.
@@ 2.2 Efficient Embedding Systems
Systems like Sentence Transformers [10], Txtai [8], and Weaviate [14] optimize around transformer inference rather than eliminating it. FastText [1] provides subword embeddings but lacks contextual understanding, while traditional approaches like GloVe [17] and Word2Vec [18] offer static embeddings without modern contextual awareness. Our methodology occupies a novel position by completely bypassing transformer inference while maintaining competitive semantic performance through optimized aggregation of pre-trained token representations. Recent work on quantization [12, 15] and efficient retrieval [35, 29] has addressed specific aspects of efficiency, but none achieve the comprehensive latency reductions demonstrated in our approach.
@@ 3.2 Mathematical Formulation
Batch processing extends this through concatenated token sequences and parallel aggregation, maintaining linear complexity while maximizing hardware utilization. The formulation draws inspiration from traditional static embeddings [17, 18] while incorporating modern optimization techniques.
@@ 3.3 System Architecture and Optimizations
The Rust implementation [6] with Axum framework [2] achieves 8% higher throughput through superior hyper/tokio integration. Key optimizations include static embedding lookup via single tensor operations, SIMD JSON processing with 256-bit vector instructions, memory prefetching reducing cache misses by 30-50%, and zero-copy binary serialization eliminating memory copies. The Candle framework [4] provides efficient tensor operations and automatic backend selection.
Three response formats cater to different needs: JSON for compatibility, binary IEEE754 for maximum performance, and JSONL for streaming scenarios. The architecture incorporates lessons from efficient inference systems [34] and high-performance computing principles [29].
@@ 4.1 Experimental Setup
Our evaluation employs wrk with specialized Lua scripts for single text processing, batch processing, JSON Lines streaming, and variable batch sizes. Configuration uses 12 threads with 400 concurrent connections over 30-second test durations. We evaluate on 8 representative MTEB tasks [19] covering classification, clustering, retrieval, similarity, and pair classification, comparing against Sentence-BERT, GTE-tiny, BGE-micro, TensorRT-optimized transformers, and traditional static embeddings. Additional evaluation includes BEIR [22] for zero-shot retrieval and established benchmarks [23, 24, 25] for comprehensive assessment.
@@ 6.1 Technical Design Justifications
Rust provides native zero-copy support, compile-time memory safety, and excellent SIMD support, achieving 2.8 million requests per second in asynchronous performance. The language selection enables predictable latency without GC pauses and direct access to CPU vector instructions, drawing from modern systems programming principles [6].
@@ 6.2 Application Scope and Limitations
However, semantic degradation of 5-15% compared to contextual models represents a fundamental trade-off, with limitations in contextual disambiguation, multilingual performance achieving only 17-23% of English effectiveness, and domain specificity requiring careful application selection. These limitations align with known challenges in static embedding approaches [1, 17] while providing unprecedented latency benefits.
@@ 7 Conclusion
The 20× efficiency improvement over traditional approaches enables new real-time applications while reducing computational requirements. Future work includes hybrid contextualization frameworks [33], adaptive quantization techniques [12, 15], and cross-modal static embeddings to further bridge computational efficiency and semantic precision.
@@
We acknowledge evaluation on 8 representative MTEB tasks rather than the full benchmark, performance comparisons based on published results, theoretical ablation studies, proprietary implementation, and primary optimization for English with multilingual limitations. The benchmarking infrastructure [16] enables verification of performance characteristics while the core implementation remains proprietary.
@@ Acknowledgments
We thank the developers of open-source tools including Rust [6], Axum [2], Candle [4], and the broader open-source community for enabling this research.

### v2

@@ Abstract
keywords:
Ultra-low latency, text embeddings, Static token lookup, real-time applications, Zero-copy serialization, Edge deployment
@@ 1 Introduction
Text embeddings have become fundamental to natural language processing applications, yet transformer-based approaches introduce prohibitive latency for real-time scenarios requiring sub-millisecond response times. While models like BERT [devlin2018bert] achieve state-of-the-art semantic quality, their multi-layer attention mechanisms create computational bottlenecks that limit deployment in latency-sensitive environments.
@@ 2.1 Transformer-Based Embedding Models
Transformer architectures from BERT [devlin2018bert] to Sentence-BERT [reimers2019sentence] have established contemporary standards for semantic representations through multi-layer attention mechanisms [vaswani2017attention]. While achieving superior quality, their computational complexity renders them unsuitable for ultra-low latency scenarios. Compression techniques like DistilBERT [sanh2019distilbert] and TinyBERT [jiao2019tinybert] reduce model sizes but retain full transformer inference, fundamentally limiting latency improvements. Recent advances in efficient transformers [tay2023transformer] and model compression [xiao2023survey] have explored various optimization paths while maintaining transformer architecture.
@@ 2.2 Efficient Embedding Systems
Systems like Sentence Transformers [reimers2019sentence2], Txtai [neuml2020txtai], and Weaviate [weaviate2019] optimize around transformer inference rather than eliminating it. FastText [bojanowski2017enriching] provides subword embeddings but lacks contextual understanding, while traditional approaches like GloVe [pennington2014glove] and Word2Vec [mikolov2013efficient] offer static embeddings without modern contextual awareness. Our methodology occupies a novel position by completely bypassing transformer inference while maintaining competitive semantic performance through optimized aggregation of pre-trained token representations. Recent work on quantization [shen2020qbert, zafrir2019q8bert] and efficient retrieval [khattab2020colbert, johnson2019billion] has addressed specific aspects of efficiency, but none achieve the comprehensive latency reductions demonstrated in our approach.
@@ 3.2 Mathematical Formulation
Batch processing extends this through concatenated token sequences and parallel aggregation, maintaining linear complexity while maximizing hardware utilization. The formulation draws inspiration from traditional static embeddings [pennington2014glove, mikolov2013efficient] while incorporating modern optimization techniques.
@@ 3.3 System Architecture and Optimizations
The Rust implementation [klabnik2018rust] with Axum framework [tokio2021axum] achieves 8% higher throughput through superior hyper/tokio integration. Key optimizations include static embedding lookup via single tensor operations, SIMD JSON processing with 256-bit vector instructions, memory prefetching reducing cache misses by 30-50%, and zero-copy binary serialization eliminating memory copies. The Candle framework [huggingface2023candle] provides efficient tensor operations and automatic backend selection.
Three response formats cater to different needs: JSON for compatibility, binary IEEE754 for maximum performance, and JSONL for streaming scenarios. The architecture incorporates lessons from efficient inference systems [lee2019latency] and high-performance computing principles [johnson2019billion].
@@ 4.1 Experimental Setup
Our evaluation employs wrk with specialized Lua scripts for single text processing, batch processing, JSON Lines streaming, and variable batch sizes. Configuration uses 12 threads with 400 concurrent connections over 30-second test durations. We evaluate on 8 representative MTEB tasks [muennighoff2022mtEB] covering classification, clustering, retrieval, similarity, and pair classification, comparing against Sentence-BERT, GTE-tiny, BGE-micro, TensorRT-optimized transformers, and traditional static embeddings. Additional evaluation includes BEIR [thakur2021beir] for zero-shot retrieval and established benchmarks [wang2018glue, rajpurkar2016squad, conneau2018xnli] for comprehensive assessment.
@@ 6.1 Technical Design Justifications
Rust provides native zero-copy support, compile-time memory safety, and excellent SIMD support, achieving 2.8 million requests per second in asynchronous performance. The language selection enables predictable latency without GC pauses and direct access to CPU vector instructions, drawing from modern systems programming principles [klabnik2018rust].
@@ 6.2 Application Scope and Limitations
However, semantic degradation of 5-15% compared to contextual models represents a fundamental trade-off, with limitations in contextual disambiguation, multilingual performance achieving only 17-23% of English effectiveness, and domain specificity requiring careful application selection. These limitations align with known challenges in static embedding approaches [bojanowski2017enriching, pennington2014glove] while providing unprecedented latency benefits.
@@ 7 Conclusion
The 20× efficiency improvement over traditional approaches enables new real-time applications while reducing computational requirements. Future work includes hybrid contextualization frameworks [lewis2020retrieval], adaptive quantization techniques [shen2020qbert, zafrir2019q8bert], and cross-modal static embeddings to further bridge computational efficiency and semantic precision.
@@
We acknowledge evaluation on 8 representative MTEB tasks rather than the full benchmark, performance comparisons based on published results, theoretical ablation studies, proprietary implementation, and primary optimization for English with multilingual limitations. The benchmarking infrastructure [lansiaux2025benchmarks] enables verification of performance characteristics while the core implementation remains proprietary.
@@ Acknowledgments
We thank the developers of open-source tools including Rust [klabnik2018rust], Axum [tokio2021axum], Candle [huggingface2023candle], and the broader open-source community for enabling this research.
Declarations
7.1 Funding
Not applicable
7.2 Conflict of interest/Competing interests
Not applicable